OpenStudy (anonymous):

Identify whether the series summation of 15 open parentheses 4 close parentheses to the I minus 1 power from 1 to infinity is a convergent or divergent geometric series and find the sum, if possible.

OpenStudy (anonymous):

\[15\sum_{k=1}^{\infty}4^{-k}\]?

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

well \[\sum_{k=0}^{\infty}\frac{1}{4^k}=\frac{1}{1-\frac{1}{4}}=\frac{4}{3}\] so this would in fact converge. then multiply by 15

OpenStudy (anonymous):

i always have to think "do you start at one or at zero" seems like i could remember...

OpenStudy (anonymous):

wouldn't 4 ^(-infinity) converge to 0?